### v1

# SEMPO - Retrieving poles, residues and zeros in the complex frequency plane from an arbitrary spectral response [^1] This work was funded by the French National Research Agency ANR Project DILEMMA (ANR-20-CE09-0027)

Abstract
The Singularity Expansion Method Parameter Optimizer - SEMPO - is a toolbox to extract the complex poles, zeros and residues of an arbitrary response function acquired along the real frequency axis. SEMPO allows to determine this full set of complex parameters of linear physical systems from their spectral responses only, without prior information about the system. The method leverages on the Singularity Expansion Method of the physical signal. This analytical expansion of the meromorphic function in the complex frequency plane motivates the use of the Cauchy method and auto-differentiation-based optimization approach to retrieve the complex poles, zeros and residues from the knowledge of the spectrum over a finite and real spectral range. Both approaches can be sequentially associated to provide highly accurate reconstructions of physical signals in large spectral windows. The performances of SEMPO are assessed and analysed in several configurations that include the dielectric permittivity of materials and the optical response spectra of various optical metasurfaces.
keywords:
Auto-differentiation, Cauchy method, complex analysis, poles, residues, singularity expansion method, zeros
[…]
[1]Aix Marseille Univ, CNRS, Centrale Mediterranee, Institut Fresnel, 13013 Marseille, France
1 Introduction
The spectral response of a linear physical system carries rich information on the system and on its interaction with excitation fields. The complex frequency space is well suited to perform this analysis, and the poles of the different transfer operators or scattering channels provide much of this information on the system. Different methods have been developed in order to model the response of the system to a given excitation with parameters determined in the complex frequency space including the Singularity Expansion Method (SEM), the Resonant State Expansion, or the quasi-normal modes theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
The SEM provides an expansion of any physical signal, such as the elements of a scattering operator, in terms of its poles and residues or alternatively in terms of its poles and zeros determined in the complex frequency plane. The expressions resulting from the distributions of log-singularities (poles and zeros) and residues can be used to approximate functions at any frequency over large spectral windows, with small and discrete sets of parameters [11, 12, 13, 14, 15]. They provide simple models to describe the behaviour of scatterers and cavities [16, 4, 17, 18]. The behaviour of physical systems and their interaction with excitation signals can be explained in terms of poles and zeros [19, 20, 21, 22, 23, 24].
Information about the distribution in the complex frequency plane being encoded in the spectral response of the physical system, the challenge is therefore to decode this information. Different methods have been proposed to retrieve the log-singularities and residues from a spectral response either obtained in the complex $\omega$ plane or along the real frequency axis. When the available data are acquired or generated in the complex frequency plane, the calculation of the poles and residues (or equivalently the poles and zeros) is quite straightforward through contour integrals [25, 26, 22, 27]. However, in practice, many spectra are acquired over a real spectral range $[\omega_{A},\omega_{B}]\subset\mathbb{R}$. In the latter case, much fewer methods exist but this longstanding problem has been attracting more and more attention [28, 29, 30, 31, 32]. Existing methods mainly approximate the response functions as meromorphic functions, the parameters of which are either the poles and zeros (Cauchy method [31] and AAA algorithm [32] indirectly), or the poles and residues (matrix pencil method [28], harmonic inversion [29], vector fitting [30]). They rely on matrix algebra, but can be quite sensitive to an initial guess of some parameters, as in the case of the vector fitting method.
Here, we describe two approaches to identify, in the complex frequency plane, the log-singularities (poles and zeros) and residues of a physical system from spectra acquired over a finite range of real frequencies. The first approach is based on the Cauchy method [14, 31, 33] and works by expressing the function through a rational approximation. In this work, we describe the working principle of the original method and explain how it can be improved and optimized using the properties of physical systems. The second method relies on auto-differentiation [34, 35, 36, 37, 38, 24] to solve optimization problems through gradient-descent. The formulation of these optimization problems is detailed, as well as their application to retrieve the SEM.
The physical signal that we consider is a transfer function such as an element of the scattering $S$-matrix that is represented by a meromorphic function $h(\omega)$ which can be expanded using the SEM [1, 39, 4, 40, 5], i.e.
$h(\omega)=h_{NR}+\sum_{\ell}\frac{r^{(\ell)}}{\omega-p^{(\ell)}.}$ (1)
$h_{NR}$ is the non-resonant term, $\{p^{(\ell)}\}_{\ell}$ is the set of simple poles of $h(\omega)$ (we assume that the singularities are all poles of order $1$), and $r^{(\ell)}$ is the residue associated with the pole $p^{(\ell)}$. Equivalently, $h(\omega)$ can be expressed through the Singularity and Zero Factorization (SZF) [2, 41, 3, 5]:
$h(\omega)=\eta_{0}~{}\frac{\displaystyle\prod_{\ell}(\omega-z^{(\ell)})}{\displaystyle\prod_{\ell}(\omega-p^{(\ell)})}$ (2)
with $z^{(\ell)}$ the zeros of $h(\omega)$, and $\eta_{0}$ a known constant which depends on $z^{(\ell)}$, $p^{(\ell)}$, and the response $h(a)$ to an arbitrary frequency $a$.
The distribution of poles $p^{(\ell)}$, zeros $z^{(\ell)}$, and residues $r^{(\ell)}$ fully characterize the response function $h(\omega)$. Therefore, the log-singularities are intrinsic properties of physical systems, and their retrieval is relevant. However, there is a potentially infinite number of parameters, which implies that they are never all taken into account since the SZF and SEM must be numerically truncated. In the following we describe how to retrieve the poles, residues and zeros once we perform this truncature. Furthermore, we show how to combine both methods to improve the overall performance.
2 The Cauchy method for rational approximations
2.1 Overview of the original method
Let us first present an overview of the Cauchy method to retrieve the poles and zeros of a function from a set of measurements acquired along a real finite spectral range.
We consider Equation (1) truncated at the $M_{p}^{\mathrm{th}}$ order for the poles, or Equation (2) also truncated at the $M_{z}^{\mathrm{th}}$ order for the zeros. Since the response function $h(\omega)$ is meromorphic, it can also be written as a rational function:
$h(\omega)=\frac{f(\omega)}{g(\omega)}$ (3)
with $g(\omega)$ a polynomial of degree $M_{p}$, and $f(\omega)$ a polynomial of degree $M_{z}$. The degree $M_{p}$ of $g(\omega)$ is initialized to the number of expected poles, and is case dependent. The roots of $g(\omega)$ and $f(\omega)$ are respectively the poles and zeros of $h(\omega)$. The calculation of the poles and zeros is thus tantamount to the calculation of the coefficients $\mathbf{a}=(a_{0},a_{1},...,a_{M_{z}})^{T}$ of $f(\omega)$ and $\mathbf{b}=(b_{0},b_{1},...,b_{M_{p}})^{T}$ of $g(\omega)$. For any frequency $\omega$, we have
$f(\omega)=\mathbf{x}^{(z)}\mathbf{a}$ (4)
and
$g(\omega)=\mathbf{x}^{(p)}\mathbf{b}$ (5)
with $\mathbf{x}^{(z)}=(1,\omega,\omega^{2},...,\omega^{M_{z}})$, $\mathbf{x}^{(p)}=(1,\omega,\omega^{2},...,\omega^{M_{p}})$. Equation (3) can thus be recast as
$\left[\mathbf{x}^{(z)},~{}-h(\omega)\mathbf{x}^{(p)}\right]\begin{bmatrix}\mathbf{a}\\ \mathbf{b}\\ \end{bmatrix}=0$ (6)
The data are acquired at $N$ frequencies $\omega_{n}$, $1\leq n\leq N$, between $\omega_{inf}$ and $\omega_{sup}\in\mathbb{R}$:
$\omega_{inf}\leq\omega_{1}<\omega_{2}<...<\omega_{N}\leq\omega_{sup}$ (7)
Let us write $h(\omega_{n})=h_{n}$ the value of the response function at $\omega_{n}$. The following matrix notation of Equation (6) is obtained:
$[\mathbf{A},-\mathbf{B}]\begin{bmatrix}\mathbf{a}\\ \mathbf{b}\\ \end{bmatrix}=\mathbf{0}.$ (8)
The matrix $\mathbf{A}$ reads as
$\mathbf{A}=\begin{bmatrix}1&\omega_{1}&...&\omega_{1}^{M_{z}}\\ 1&\omega_{2}&...&\omega_{2}^{M_{z}}\\ &...&\\ 1&\omega_{N}&...&\omega_{N}^{M_{z}}\\ \end{bmatrix}\$ (9)
and $\mathbf{B}$ as
$\mathbf{B}=\begin{bmatrix}h_{1}&h_{1}\omega_{1}&...&h_{1}\omega_{1}^{M_{p}}\\ h_{2}&h_{2}\omega_{2}&...&h_{2}\omega_{2}^{M_{p}}\\ &...&\\ h_{N}&h_{N}\omega_{N}&...&h_{N}\omega_{N}^{M_{p}}\\ \end{bmatrix}.$ (10)
The Cauchy method constrains the kernel of the matrix $\mathbf{C}=[\mathbf{A},~{}-\mathbf{B}]$ in order to obtain a unique algebraic solution up to a scaling factor. The number of unique frequencies $N$ is usually greater than the number of columns $2+M_{p}+M_{z}$ of $\mathbf{C}$. It follows that the rank $r$ of $\mathbf{C}$ is given by
$r=M_{z}+M_{p}+2-K$ (11)
with $K$ the dimension of its kernel. The Cauchy method aims at setting the dimension of the kernel to $K=1$.
Figure 1: (a) Si nanodisks of diameter $d=160$ nm and height $h=130$ nm, arranged in a 2D array of period $p=420$ nm over a substrate of glass. The signal $\mathbf{h}$ is the $0^{th}$-order optical reflection coefficient at normal incidence in air. (b) Relative singular values of the matrix $\mathbf{C}=[\mathbf{A}~{}-\mathbf{B}]$ with $\mathbf{A}$ and $\mathbf{B}$ defined in Equations (9) and (10). The singular values are sorted by decreasing value, and divided by the first one, $\sigma_{max}$. In the classical Cauchy method, the rank $r$ is set to the index of the smallest singular value after which no significant variation of the ratio $\sigma_{i}/\sigma_{max}$ is observed. (c) Reflection coefficient $h(\omega)=\tilde{r}_{00}(\omega)$ retrieved via the classical Cauchy method (red curve), and compared to simulated data (blue markers). The absolute error $|h_{i}-\hat{h}_{i}|$, with $h_{i}$ the target and $\hat{h}_{i}$ the reconstructed value, is plotted (black curve), and shows the high accuracy of the method, with a relative $L_{2}$ error $e^{(2)}_{22,11,10}=4.55\times 10^{-1}$%. (d) Distribution of poles and zeros retrieved by the Cauchy method in the complex $\omega$ plane. Some poles and zeros are located outside the spectral window of interest.
The Cauchy method first requires an estimation of the rank $r$. This is achieved by first choosing large values of $M_{p}$ and $M_{z}$, and then performing a Singular Value Decomposition (SVD) of $\mathbf{C}$. The rank $r$ corresponds to the index of the smallest singular value not considered as null, when all the singular values are sorted in descending order.
We illustrate this process in an example. The response function $h(\omega)$ is the $0^{th}$ order reflection coefficient $h(\omega)=r_{00}(\omega)$ of a 2D array of Si nanodisks illustrated in Figure 1 (a). The initial numbers of poles and zeros are $M_{p}=20$ and $M_{z}=19$. The first null singular value is identified by the inflexion point in Figure 1 (b), which displays the sorted relative singular values.
Once the rank $r$ is estimated, we can select $M_{z}$ and $M_{p}$ in order to have $K=1$ in Equation (11). Let us point out that, classically, $M_{z}$ is equal to $M_{p}-1$ in the Cauchy method. After these values of $M_{z}$ and $M_{p}$ have been selected, we can calculate one solution to Equation (8) by first performing a QR decomposition of $\mathbf{A}$:
$\mathbf{A}=\mathbf{Q}\begin{bmatrix}\mathbf{R}_{11}\\ 0\\ \end{bmatrix}$ (12)
with $\mathbf{Q}$ a $N\times N$ orthogonal matrix and $\mathbf{R}_{11}$ a full-rank $(M_{z}+1)\times(M_{z}+1)$ upper triangular matrix. By applying $\mathbf{Q}^{T}$ to $\mathbf{C}$, we obtain
$\begin{bmatrix}\mathbf{R}_{11}&\mathbf{R}_{12}\\ 0&\mathbf{R}_{22}\\ \end{bmatrix}\begin{bmatrix}\mathbf{a}\\ \mathbf{b}\\ \end{bmatrix}=0$ (13)
In Equation (13), $\mathbf{b}$ must satisfy
$\mathbf{R}_{22}\mathbf{b}=\mathbf{0}.$ (14)
The vector $\mathbf{b}$ is thus set to the eigenvector associated with the 00 eigenvalue of $\mathbf{R}_{22}$, which is uniquely defined up to a multiplicative constant since $K=1$. We can then obtain $\mathbf{a}$ from Equation (13):
$\mathbf{a}=-\mathbf{R}_{11}^{-1}\mathbf{R}_{12}\mathbf{b}.$ (15)
Knowing the polynomial coefficients $\mathbf{a}$ and $\mathbf{b}$, $f(\omega)$ and $g(\omega)$ can be calculated, and the SZF expression can be retrieved. The poles $p^{(\ell)}$ are obtained by calculating the roots of $g(\omega)$, and the zeros $z^{(\ell)}$ through the roots of $f(\omega)$. The Cauchy method thus provides a mean to reconstruct a signal through the SZF.
We test the classical Cauchy method and show its accuracy in Figure 1 (c,d), using the reflection coefficient $h(\omega)=r_{00}(\omega)$ introduced in Figure 1 (a). The response function is evaluated at $30$ frequencies ranging from $2.3$ to $4.7\times 10^{15}$ rad/s. The value of the rank is $r=22$, resulting in $M_{p}=11$ and $M_{z}=M_{p}-1=10$. Despite the small numbers of poles and zeros, we observe a good fitting over a large spectral window, with a relative $L_{2}$ error $e^{(2)}_{21,11,10}$ of $4.55\times 10^{-1}$%, defined as
$e^{(2)}_{r,M_{p},M_{z}}\left[\hat{\mathbf{h}},\mathbf{h}\right]=\frac{{||\hat{\mathbf{h}}-\mathbf{h}||}_{2}}{{||\mathbf{h}||_{2}}}$ (16)
where $\mathbf{h}=\left(h_{1},...,h_{N}\right)$ is the experimental data vector, and $\hat{\mathbf{h}}=\left(\hat{h}_{1},...,\hat{h}_{N}\right)$ is the reconstructed response vector using the parameters $r$, $M_{p}$ and $M_{z}$. The norm of a vector $\mathbf{v}=\left(v_{1},...,v_{N}\right)$ reads as
${||\mathbf{v}||}_{2}=\sqrt{\sum_{i=1}^{N}|v_{i}|^{2}}.$ (17)
2.2 Accuracy-driven Cauchy method
While the classical Cauchy method provides good results in most situations, it can be improved by modifying some steps, in particular the calculation of the rank and the determination of the final numbers of poles and zeros. In the classical method, the rank is estimated by performing an SVD on $\mathbf{C}$, and is given by the index of the last non-null singular value, as previously shown in Figure 1 (b). Consequently, there is no guarantee that the dimension $K$ of the kernel will be equal to $1$. In addition, there is usually no a priori information imposing $M_{z}=M_{p}-1$. The accuracy of the Cauchy method can thus be easily enhanced by enforcing a kernel dimension $K$ of $1$, and by sweeping through different values of $M_{p}-M_{z}\geq 0$.
We first choose a large initial number $M_{p,0}$ of poles, and set the initial number of zeros $M_{z,0}$ to that same value. This results in a matrix $\mathbf{C}_{0}=[\mathbf{A}_{0},-\mathbf{B}_{0}]$ of dimension $N\times 2(M_{p,0}+1)$, with $\mathbf{A}_{0}$ and $\mathbf{B}_{0}$ defined in Equations (9) and  (10). We then perform an SVD of $\mathbf{C}_{0}$, and use the same method as described in Figure 1 (b) to obtain a value $r_{\max}$ instead of an approximation of the rank this time. If $r_{\max}$ is even, we add $1$ to get an odd number.
This leads to two new maximum numbers of poles and zeros $M_{p,\max}=M_{z,\max}=r_{\max}/2$. They are used as boundaries as we iteratively increase the numbers of zeros $M_{z}$ and poles $M_{p}$:
$1\leq M_{z}\leq M_{z,\max}$ (18)
$M_{z}+\Delta D\leq M_{p}\leq M_{p,\max}.$ (19)
$\Delta D$ is a free parameter of the method corresponding to the maximum difference between the number of poles and the number of zeros. A total of $\Lambda=M_{z,\max}\times(M_{z,\max}-\Delta D)$ couples $(M_{z},M_{p})$ are tested. These combinations are identified by the index $\lambda$, i.e. $1\leq\lambda\leq\Lambda$.
For each value of $\lambda$, a matrix $\mathbf{C}_{\lambda}$ is calculated, defined once again through Equations (9) and  (10). We can then perform another SVD:
$\mathbf{C}_{\lambda}=\mathbf{U}\begin{bmatrix}\mathrm{\mathbf{diag}(\sigma_{1},...,\sigma_{r+1})}\\ 0\\ \end{bmatrix}\mathbf{V}^{H}$ (20)
where $\mathbf{U}$ and $\mathbf{V}$ are two unitary matrices, $\mathbf{V}^{H}$ is the conjugate transpose of $\mathbf{V}$, $r=M_{z}+M_{p}+2$, and $\sigma_{1},\sigma_{2},...,\sigma_{r_{+}1}$ are the singular values of $\mathbf{C}_{\lambda}$. By setting $\sigma_{r+1}$ to 00, we enforce $r$ as the rank of the resulting matrix $\hat{\mathbf{C}}_{\lambda}$, ensuring a dimension of the kernel $K=1$. We then perform the classical Cauchy method on all matrices $\hat{\mathbf{C}}_{\lambda}$, resulting in $\Lambda$ solutions $[\mathbf{a}_{\lambda},\mathbf{b}_{\lambda}]$, each associated with a reconstruction $\hat{\mathbf{h}}_{\lambda}$ using Equations (3) to  (5).
Figure 2: Reconstruction of the relative permittivity $\varepsilon_{\mathrm{Au}}(\omega_{i})=h(\omega)$ of $\mathrm{Au}$, using the classical Cauchy method (dashed, red curve) and the ADC method (blue curve). A better reconstruction is obtained with the ADC method, which yields a relative $L_{2}$ error $e^{(2)}_{18,9,8}=2.53\times 10^{-3}$%, lower than that obtained via the classical Cauchy method, which is $e^{(2)}_{20,10,9}=5.67\times 10^{-3}$%.
The relative $L_{2}$ error $e^{(2)}_{r,M_{p},M_{z}}\left[\mathbf{h}_{\lambda},\mathbf{h}\right]$ corresponding to each solution is evaluated. The optimal solution $\mathbf{h}_{\lambda^{*}}$ is chosen, defined as
$\lambda^{*}=\operatorname*{arg\,min}_{\lambda}e^{(2)}_{r,M_{p},M_{z}}\left[\mathbf{h}_{\lambda},\mathbf{h}\right].$ (21)
We refer to this enhanced approach as the Accuracy-Driven Cauchy method (ADC method). The ADC method is, by construction, always at least as accurate as the classical method when the same initial amount of poles $M_{p,0}$ is chosen. We compare the two approaches in Figure 2, where the response function is the relative permittivity of gold $h(\omega)=\varepsilon_{\mathrm{Au}}(\omega)$, and assess the increased performance of the ADC method compared to the conventional Cauchy method.
2.3 Physics-informed corrections
So far, we have focused on the quality of the reconstruction without paying attention to the distributions of poles and zeros. Nevertheless, these complex frequencies are constrained in the case of physical systems and signals. We propose a means to take into account the properties of physical systems while maintaining an accurate fitting of the data.
If $h(\omega)$ is the response or the transfer function of a physical system, it is Hermitian-symmetric, i.e. its poles and zeros come in pairs $(p^{(\ell)},-\overline{p^{(\ell)}})$ and $(z^{(\ell)},-\overline{z^{(\ell)}})$, and the constant $\eta_{0}$ is real-valued. We can rewrite the truncated SZF [5] as
$\frac{h(\omega)}{\eta_{0}}=\frac{\displaystyle\prod_{\ell=1}^{M_{z,\mathrm{i}\mathbb{R}}}(\omega-\mathrm{i}z_{\mathrm{i}\mathbb{R}}^{(\ell)})}{\displaystyle\prod_{\ell=1}^{M_{p,\mathrm{i}\mathbb{R}}}(\omega+\mathrm{i}p_{\mathrm{i}\mathbb{R}}^{(\ell)})}\frac{\displaystyle\prod_{\ell=1}^{M_{z,\mathbb{C}}}\left[(\omega-z_{\mathbb{C}}^{(\ell)})(\omega+\overline{z_{\mathbb{C}}^{(\ell)}})\right]}{\displaystyle\prod_{\ell=1}^{M_{p,\mathbb{C}}}\left[(\omega-p_{\mathbb{C}}^{(\ell)})(\omega+\overline{p_{\mathbb{C}}^{(\ell)}})\right]}$ (22)
where $\mathrm{i}z_{\mathrm{i}\mathbb{R}}$ are purely imaginary zeros, $-\mathrm{i}p_{\mathrm{i}\mathbb{R}}$ are purely imaginary poles, $z_{\mathbb{C}}$ are non-imaginary zeros, and $p_{\mathbb{C}}$ are non-imaginary poles. The number of poles and zeros read as $M_{z}=M_{z,\mathrm{i}\mathbb{R}}+M_{z,\mathbb{C}}$ and $M_{p}=M_{p,\mathrm{i}\mathbb{R}}+M_{p,\mathbb{C}}$.
In order to take into account this Hermitian symmetry, we add negative frequencies, effectively multiplying the number of frequencies by 2. For every frequency $\omega_{i}$ associated with the value $h_{i}$, the opposite frequency $-\omega_{i}$ is added, along with $\overline{h_{i}}$. This operation is performed at the beginning of the Cauchy method, right before the step corresponding to Equation (6).
The resulting distributions of poles and zeros are not perfectly Hermitian-symmetric, whether we use the ADC method or the classical Cauchy method. We correct for discrepancies by removing all the poles and zeros with a negative real part, before replacing them with the Hermitian-symmetrics of the poles and zeros with a positive real part. We thus explicitly impose the symmetry expressed in Equation (22) to reconstruct $h(\omega)$. In the ADC method, the poles and zeros are replaced before calculating the errors $e^{(2)}_{r,M_{p},M_{z}}[\mathbf{h}_{\lambda},\mathbf{h}]$ in order to ensure that the optimal Hermitian-symmetric solution is chosen among all the calculated options.
In addition, the poles and zeros located too far from the window of interest act as offsets and have little influence over the variations of $h(\omega)$ at real frequencies. We thus remove them from the sets of poles and zeros, but modify the constant term $\eta_{0}$ in order to take into account their contributions:
$\eta_{0}^{\prime}=\eta_{0}\frac{\displaystyle\prod_{z_{f}}\left(-|z_{f}|^{2}\right)}{\displaystyle\prod_{p_{f}}\left(-|p_{f}|^{2}\right)}$ (23)
where $\eta_{0}^{\prime}$ is the modified constant term, and $z_{f}$ and $p_{f}$ are the zeros and poles far from the spectral range of interest. They are considered as such when their distance to the origin exceeds a certain threshold $\rho$, which is by default set to $5$ times the spectral range. The calculation of $\eta_{0}^{\prime}$ is also performed before the calculation of the errors in the ADC method.
Finally, we propose to further modify the distribution of poles to account for the usually assumed stability of physical systems or signals. Stable poles possess a negative imaginary part. Therefore, we penalize the solutions associated with unstable poles in the ADC method by multiplying the errors $e^{(2)}_{r,M_{p},M_{z}}\left[\mathbf{h}_{\lambda},\mathbf{h}\right]$ by $(1+M_{inst,p})$ with $M_{inst,p}$ the number of poles with a strictly positive imaginary part. What is more, we add a negative imaginary part $-\mathrm{i}q_{0}$ to the poles located on the real frequency axis once the optimal solution is obtained. $q_{0}$ is yet another free parameter set to $10^{-5}$ by default. Since the effect of this modification on the zeros is not easily obtained, this last step is performed after converting the SZF into the SEM, i.e. by considering the poles and their associated residues instead of considering the poles and zeros.
The residue $r^{(m)}$ associated with a pole $p^{(m)}$ is calculated as
$r^{(m)}=\eta_{0}~{}\frac{\displaystyle\prod_{\ell=1}^{M_{z}}(p^{(m)}-z^{(\ell)})}{\displaystyle\prod_{\ell\neq m}(p^{(m)}-p^{(\ell)})}$ (24)
and the non-resonant term is obtained through the expression [18]
$h_{NR}=h(\omega_{0})+\sum_{\ell=1}^{M_{p}}\frac{r^{(\ell)}}{p^{(\ell)}}$ (25)
where $\omega_{0}$ is an arbitrary complex frequency different from the poles. In addition, we remove the poles associated with residues which have a small modulus compared to the maximum modulus. The default threshold is $1\%$ of the maximum.
Figure 3: (a) 1D gold grating over a substrate of gold, with height $h=60$ nm, width $w=100$ nm, and period $p=200$ nm. The studied function is the $0^{th}$ order reflection coefficient $r_{0}(\omega)=h(\omega)$ at normal incidence in air. (b) Approximation of $h(\omega)$, using the ADC method (blue curve, ADC) or the physics-informed version (red curve, Herm. sym. ADC). (c) Distributions of poles in the complex $\omega$ plane using the two approaches. While both versions provide similar results, i.e. an accurate fitting of the experimental data (black markers), the physics-informed method fulfils the Hermitian symmetry and requires fewer pairs of poles and zeros.
We compare the ADC method to its Hermitian-symmetric counterpart in the case of the reflection coefficient of a 1D gold grating described in Figure 3 (a). We show that the reconstruction is only slightly affected by the proposed changes in Figure 3 (b, c).
The Cauchy method may fail to provide accurate models in the case of rapidly varying curves, or when stability is mandated, as it will be shown in the last section of this manuscript. Despite its rapidity of execution, it must therefore be replaced in these cases by an alternative method, which is offered here by the auto-differentiation approach.
3 Auto-differentiation approach for the singularity expansion
3.1 Formulating the optimization problems
We now move onto the second approach which aims at exploiting auto-differentiation (as proposed in the PyTorch library) to obtain accurate approximations through optimization problems. Although these approaches could be applied to optimize the SZF, we focus on the retrieval of the different terms in the SEM, or in other words, the poles $p^{(\ell)}$, the residues $r^{(\ell)}$, and the non-resonant term $h_{NR}$. Similarly to the previous section, we consider the truncated SEM expression, with $M$ complex pairs of poles or purely imaginary poles:
$h(\omega)=h_{NR}+\sum_{\ell=1}^{M}\left[\frac{r^{(\ell)}}{\omega-p^{(\ell)}}-\frac{\overline{r^{(\ell)}}}{\omega+\overline{p^{(\ell)}}}\right].$ (26)
While it is possible to deal with complex values in the optimization process, current auto-differentiation tools have only recently been able to deal with such numbers, and some functions are still not properly defined in $\mathbb{C}$ [42, 43]. We thus express the parameters in $\mathbb{R}^{2}$ instead of $\mathbb{C}$. We split the $M$ poles of Equation (26) into $M_{I}$ purely imaginary poles $\mathrm{i}q^{(\ell)}$ with residues $\mathrm{i}\rho^{(\ell)}$ and $M_{C}$ complex pairs of poles $p^{(\ell)}=p_{R}^{(\ell)}+\mathrm{i}p_{I}^{(\ell)}$ with residues $r^{(\ell)}=a^{(\ell)}+\mathrm{i}b^{(\ell)}$.
We obtain a form of the SEM expression that is compatible with the auto-differentiation tools [24]:
$h(\omega)=h_{NR}+\sum_{\ell=1}^{M_{I}}\frac{\rho^{(\ell)}}{q^{(\ell)}}+\sum_{\ell=1}^{M_{C}}\left[\frac{a^{(\ell)}+\mathrm{i}b^{(\ell)}}{\omega-p_{R}^{(\ell)}-\mathrm{i}p_{I}^{(\ell)}}-\frac{a^{(\ell)}-\mathrm{i}b^{(\ell)}}{\omega+p_{R}^{(\ell)}-\mathrm{i}p_{I}^{(\ell)}}\right].$ (27)
The non-resonant term is also a real-valued parameter due to the Hermitian symmetry. The explicit discrimination between the $M_{I}$ imaginary poles and the $M_{C}$ pairs of poles is not necessary, but it can help in improving the results if imaginary poles are expected. In addition, it reduces the number of parameters from $1+4M$ to $1+2M_{I}+4M_{C}$ which can slightly improve the speed of the optimization process.
We wish to find a set of parameters $\mathcal{P}$ defined as
$\mathcal{P}=\left\{h_{NR},~{}\left(\rho^{(\ell)}\right)_{\ell},\left(q^{(\ell)}\right)_{\ell}\left(a^{(\ell)}\right)_{\ell},\left(b^{(\ell)}\right)_{\ell},\left(p_{R}^{(\ell)}\right)_{\ell},\left(p_{I}^{(\ell)}\right)_{\ell}\right\}$ (28)
for which a small value of the arbitrary loss function $\mathcal{L}$ is reached. Ideally, we want to retrieve the optimal set of parameters $\mathcal{P}^{*}$ defined as
$\mathcal{P}^{*}=\operatorname*{arg\,min}_{\mathcal{P}}\mathcal{L}[\mathbf{h},\hat{\mathbf{h}},\mathcal{P}]$ (29)
where $\mathcal{L}[h_{i},\hat{h}_{i},\mathcal{P}]$ evaluates the distance of $\hat{h}_{i}$ to $h_{i}$ using the parameter set $\mathcal{P}$ in the truncated SEM from Equation (26) to reconstruct the function. The minimization of the loss function is performed via a gradient-descent(-like) method. Classically, the loss function is limited to the set of functions for which this gradient can be computed.
Nevertheless, the recent development of auto-differentiation (for machine learning in particular), has provided means to explore more exotic loss functions. The tools provided in the PyTorch library allow to freely design the loss function while the gradient is automatically calculated by the library through the chain rule.
The loss function that we use requires the definition of the relative $L_{2}$ and $L_{\infty}$ errors:
$e_{q}\left[\mathbf{h},\hat{\mathbf{h}}\right]=\frac{{||\mathbf{h}-\hat{\mathbf{h}}||}_{q}}{{||\mathbf{h}||_{q}}}$ (30)
with $q\in\{2,\infty\}$, i.e.
$\begin{gathered}{||\mathbf{v}||}_{2}=\sqrt{\sum_{i=1}^{N}|v_{i}|^{2}}\\ {||\mathbf{v}||}_{\infty}=\max_{i}|v_{i}|.\end{gathered}$ (31)
The loss function itself is then defined as
$\mathcal{L}[\mathbf{h},\hat{\mathbf{h}},\mathcal{P}]=\alpha_{1}e^{(2)}\left[\mathbf{h},\hat{\mathbf{h}}\right]+\alpha_{2}\left\lVert\frac{\mathbf{h}-\hat{\mathbf{h}}}{\mathbf{h}}\right\rVert_{\infty}+\alpha_{3}\left<\frac{\delta\mathbf{h}_{R}}{\mathbf{h}_{R}}\right>+\alpha_{4}\left<\frac{\delta\mathbf{h}_{I}}{\mathbf{h}_{I}}\right>$ (32)
where the notation <.><.>< . > corresponds to the averaging operation over all the sampled frequencies. $\delta\mathbf{h}_{R}$ is the vector of errors between the real part of the model $\hat{\mathbf{h}}$ and the experimental data $\mathbf{h}$, and $\mathbf{h}_{R}$ is the vector of the absolute values of the real part $\mathbf{h}$. It is used to change the weight given to the frequency $\omega_{i}$ in $\mathbf{\delta h_{R}}$. $\mathbf{\delta h_{I}}$ and $\mathbf{h}_{I}$ do the same for the imaginary part:
[…]
The values of the weight coefficients are usually case-dependent. The first coefficient $\alpha_{1}$ roughly corresponds to the average error between $\mathbf{h}$ and $\hat{\mathbf{h}}$. $\alpha_{2}$ can be used to give more weight to small variations in the amplitude. $\alpha_{3}$ and $\alpha_{4}$ also allow to strengthen the contribution of points associated with small variations or values, but in the real and imaginary parts respectively.
In an accuracy-driven optimization, these parameters, along with the numbers of poles $M_{I}$ and $M_{C}$, can be optimized by grid-search as to validate a criterion such as the minimization of the relative $L_{2}$ error between $h_{i}$ and $\hat{h}_{i}$.
However, this could lead to a fitting of the noise rather than fitting the signal brought about by the data. We have found that setting $\alpha_{1}=1$, and keeping the other weights small often provides good results. In the following examples, the values of $\mathbf{\alpha}$ are therefore intuited based on the expected response curve.
Figure 4: (a) 2D array of square Ag pillars of width $w=85$ nm and height $h=35$ nm, and period $p=130$ nm over a substrate of Ag. (b) The signal is the 00-th order reflection coefficient $h(\omega)=r_{00}(\omega)$ of the 2D array depicted in (a) illuminated in normal incidence reconstructed with the SEM in Equation (27) (red curve), and compared to simulated data (blue markers). The absolute error $|h_{i}-\hat{h}_{i}|$ (black curve) is less than $10^{-3}$ at all frequencies, and a relative $L_{2}$ error of $4.32\times 10^{-3}$ is obtained. (c) Distribution of poles in the complex $\omega$ plane.
Using the example of the 00-th order reflection coefficient of a 2D grating of silver described in Figure 4 (a), we show the result of the retrieval of the singularity expansion of Equation (27) in Figure 4 (b,c), with $M_{I}=0$ imaginary poles, $M_{C}=7$ complex pairs of poles, and $\mathbf{\alpha}=(1,0,0.2,0.2)$.
These weight coefficients are such that the main focus is put on the relative $L_{2}$ error, with a smaller contribution of the errors on the real and imaginary parts, individually. We chose them after observing variations of equal amplitude of the real and imaginary parts in the target function. The non-imaginary poles were initialized with their real parts evenly distributed from $2.5$ to $4.9\times 10^{15}$ rad/s, and their imaginary part set to $-0.05$ times their real parts. We obtain a relative $L_{2}$ error of $4.32\times 10^{-3}$%, with a Hermitian-symmetric solution.
3.2 Other forms of the singularity expansion
We can cast the SEM expression into different forms which highlight different properties of physical systems:
  • The classical SEM expression in Equation (1) exhibits resonant terms associated with singularities and residues in the complex $\omega$ plane, represented as peaks, and intuitively introduces the analytical continuation of $h(\omega)$ in the complex plane.
  • The Generalized Drude-Lorentz model (GDL) [44, 24] models the physical systems as sets of sub-units, each behaving as oscillators described by classical mechanics. The sub-units can be identified as the microscopic-scale particles constituting the system.
Each of these forms is truncated and written in a way that is compatible with the optimization problem, similarly to the classical SEM in Equation (26).
As shown in reference [24], we can recast the singularity expansion into the GDL via a change of variables for the $M_{C}$ symmetric terms:
$\omega_{0,\ell}=|p_{R}^{(\ell)}|$ (35)
$\Gamma_{\ell}=-2|p_{I}^{(\ell)}|$ (36)
$s_{1,\ell}\Gamma_{l}=-2b^{(\ell)}$ (37)
$s_{2,\ell}\left(\omega_{0,\ell}^{2}+\frac{\Gamma_{\ell}}{2}^{2}\right)=-2{\mathrm{Re}}\left(r_{p}^{(\ell)}~{}\left(\omega_{0,\ell}-\mathrm{i}\frac{\Gamma_{\ell}}{2}\right)\right)$ (38)
and by rewriting the $M_{I}$ resonant terms associated with imaginary poles as
$\gamma_{\ell}=-q^{(\ell)}$ (39)
$\omega_{b,\ell}^{2}=|\rho^{(\ell)}\gamma_{\ell}|$ (40)
$\gamma_{0}=r_{0}+\sum_{\ell=1}^{M}\frac{\omega_{b,\ell}^{2}}{\gamma_{\ell}}$ (41)
$\mathrm{i}\frac{r_{0}}{\omega}+\sum_{\ell=1}^{M}\frac{\rho^{(\ell)}}{\omega-\mathrm{i}q^{(\ell)}}=\mathrm{i}\frac{\gamma_{0}}{\omega}-\sum_{\ell=1}^{M}\frac{\omega_{b,\ell}^{2}}{\omega^{2}+\mathrm{i}\omega\gamma_{\ell}}$ (42)
where we assume that there is a pole 00 at the origin, associated with a residue $\mathrm{i}r_{0}$. Let us stress that the absolute value ”|.||.|| . |” is used to impose the stability of all the poles, but it can be removed to allow for unstable effective poles. The following expression of the GDL is obtained:
$h(\omega)=h_{\mathrm{NR}}+\mathrm{i}\frac{\gamma_{0}}{\omega}-\sum_{\ell=1}^{M_{I}}\frac{\omega_{b,\ell}^{2}}{\omega^{2}+\mathrm{i}\omega\gamma_{\ell}}-\sum_{\ell=1}^{M_{C}}\frac{\mathrm{i}s_{1,\ell}\omega\Gamma_{\ell}+s_{2,\ell}\left(\omega_{0,\ell}^{2}+\frac{\Gamma_{\ell}}{2}^{2}\right)}{\left(\omega^{2}-\left(\omega_{0,\ell}^{2}+\frac{\Gamma_{\ell}}{2}^{2}\right)\right)+\mathrm{i}\omega\Gamma_{\ell}}.$ (43)
As previously stated, this formulation of the SEM is particularly adapted to the representation of sub-units of a physical system as harmonic oscillators.
We apply it to the case of the lattice of gold nanodisks presented in Figure 5 (a). The studied signal is the transmission coefficient $h(\omega)=\tilde{t}(\omega)$ at normal incidence. The reconstruction shown in Figure 5 (b,c) is obtained with $M_{I}=1$ imaginary pole, $M_{C}=8$ complex pairs of poles and $\mathbf{\alpha}=(1,0,0,0)$. The weight coefficients are such that more importance is given to the real part $h(\omega)$ which shows more complex variations than the imaginary part. We obtain a relative $L_{2}$ error of $6.31\times 10^{-4}$, over a spectral range extending from $100$nm to $900$nm.
Figure 5: (a) Lattice of gold nanodisks with height $h=50$nm, diameter $d=80$nm and period $p=500$ nm, bearing on a substrate of glass. The signal is the $0^{th}$-order transmission coefficient $h(\omega)=t_{00}(\omega)$ for a normal incidence in air. (b) Reconstruction of $h(\omega)$ with the GDL in Equation (43) (red curve), and compared to simulated data (blue markers). A relative $L_{2}$ error $e^{(2)}=6.31\times 10^{-3}$ is obtained. (c) Distribution of the poles in the complex $\omega$ plane.
4 Combined approach
The ADC method and the auto-differentiation approach aim at performing the same task, namely retrieving an analytical expression of a physical signal in the form of the SEM or SZF. Nevertheless, the two methods work in very different ways, both featuring their own limitations and use cases. This makes any direct comparison of the two methods irrelevant.
Figure 6: Influence of the frequency sampling in the ADC method on the relative $L_{2}$ error. The signal is the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$ of thickness $d=260$ nm, for a TM illumination at $17\degree$ coming from the air superstrate. $N$ frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. (a) Reconstruction of $h(\omega)$ using $N=40$ (blue), $N=60$ (orange) and $N=80$ (green) frequencies, and compared to the target (black). (b) Evolution of the relative error with the number of frequencies, with $N$ going $5$ by $5$ from $20$ to $100$.
The ADC method is faster than the auto-differentiation approach. However, it can be difficult to maintain a high accuracy while opting for a Hermitian-symmetric and stable solution, as previously shown. In addition, the method does not work well in large spectral windows showing rapid variations of the signals, and is sensitive to the sampling of data. While the latter is addressed with what is known as the adaptive Cauchy method [16], the former is not easily solved for. We illustrate both problems in Figure 6, where we apply the ADC method to retrieve the poles and zeros of the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ $h(\omega)=\tilde{r}(\omega)$, of thickness $d=260$nm, with a TM illumination at $17\degree$ in air. $N$ frequencies are uniformly picked between $0.15$ and $15.6\times 10^{15}$ rad/s. We observe in Figure 6 (a) that fast dynamics such as the one around $5\times 10^{15}$ rad/s are difficult to capture, whatever the value of $N$. In Figure 6 (b) shows that the relative $L_{2}$ does not systematically lower when $N$ is increased.
The auto-differentiation approach can provide accurate solutions even in situations where the ADC method fails. In addition, it allows for an easy implementation of constraints, in particular the Hermitian-symmetry and the stability of poles. However, this efficiency is obtained at the price of a much slower computation time depending on the hyperparameters: the optimization scheme, the desired number of iterations, the learning-rate and the learning-rate updating method, the loss function weight vector $\mathbf{\alpha}$, the sampling of the frequencies $\omega_{i}$ and values $h_{i}$, the number of poles and zeros or residues, and their initial distributions.
Figure 7: Chained Cauchy method applied to the reconstruction of the reflection coefficient of a slab of $\mathrm{TiO}_{2}$ of thickness $d=260$ nm, with a TM illumination at $17\degree$ from the air superstrate. The reconstruction is carried out in four smaller spectral windows, which are indicated by the dotted black lines.
We are particularly interested in the latter. By properly selecting the initial values of the parameters in such a way that the initial approximation is close to the target function, we can reduce the converge time towards a suitable solution, lowering the relative $L_{2}$ error, and we can often reduce the total number of poles.
We propose to obtain this initial distribution of poles and zeros through a chained version of the ADC method, which successively fits subsets of the data in smaller frequency windows. In Figure 7, we show how this applies to $h(\omega)$. We use four sub-windows, each bounded by the vertical dotted lines in Figure 7. In this case, each window is chosen so that it has the same number of frequency points, i.e. $75$ points. Each sub-window is associated with its own sets of poles and zeros and constant $\eta_{0}$.
Let us stress that despite the accurate reconstruction obtained for each sub-window, this chained ADC method does not provide a suitable solution on account of three issues:
  1. (i)
    The high number of poles and zeros associated with each sub-window prevents us from discriminating between a fitting of the curve itself and a fitting of the noise. Having too many poles and zeros generally leads to a number of degrees of freedom so high that any curve can be approximated.
  2. (ii)
    While the Hermitian-symmetry can be imposed, it is not the case for the stability criterion. Therefore, the solution is physically irrelevant.
  3. (iii)
    It is not possible to obtain a solution for the whole spectral range directly from the solutions of the sub-windows. We have thus turned one fitting problem into several ones, thereby increasing the amount of required parameters to describe the full response function.
These different problems motivate the use of another method, i.e. the auto-differentiation approach, on top of the chained ADC method to obtain a stable physical solution with fewer parameters, among which the relevant ones can be more easily identified.
Figure 8: (a) Reflection coefficient $h(\omega)=\tilde{r}(\omega)$ of a slab of $\mathrm{TiO}_{2}$ of thickness $d=260$ nm, illuminated from the air superstrate with an incidence of $17\degree$, in the TM polarization. It is reconstructed using either the SEM (blue curve), or the combined ADC and SEM approach (red curve), and compared to simulated data (black markers). A relative $L_{2}$ error $e^{(2)}=5.73\times 10^{-2}$ is obtained with the SEM approach, and $e^{(2)}=1.08\times 10^{-2}$ for the combined approach. The latter thus performs better, and is obtained in half of the time needed for the former. (b) Distribution of the poles in the complex $\omega$ plane, for the SEM (blue markers) and the combined approach (red markers). Fewer singularities are required with the ADC+SEM approach to achieve a better reconstruction than with the SEM.
For each sub-window, we convert the set of poles and zeros into a set of poles and residues, as well as a constant $h_{NR}$. A contribution weight $q_{w,\ell}$ is defined for all the pole $p^{(\ell)}$, and can be used to filter out poles that do not significantly help in the reconstruction in a sub-window. To do so, we first introduce a score $\rho_{w,\ell}$ which quantifies the variations brought about by the resonant term $h_{\ell}(\omega)$ associated with $p^{(\ell)}$:
$h_{\ell}(\omega)=\frac{r^{(\ell)}}{\omega-p^{(\ell)}}-\frac{\overline{r^{(\ell)}}}{\omega+\overline{p^{(\ell)}}}$ (44)
$d_{m}=\min_{i}|h_{p}(\omega_{i})|$ (45)
$d_{M}=\max_{i}|h_{p}(\omega_{i})|$ (46)
$\rho_{w,\ell}=1-\frac{d_{m}}{d_{M}}.$ (47)
We also introduce a score $\eta_{w,\ell}$ which quantifies the contribution of the resonant term $h_{\ell}(\omega)$ to the response function $h(\omega)$ in the sub-window:
$\eta_{w,\ell}=\frac{\sum_{i}|h(\omega_{i})-h_{\ell}(\omega_{i})|}{\sum_{i}|h(\omega_{i})|}.$ (48)
The weight $q_{w,\ell}$ is subsequently defined as
$q_{w,\ell}=\sqrt{\rho_{w,\ell}^{2}+\eta_{w,\ell}^{2}}.$ (49)
In every sub-window, the poles with a weight $q_{w,\ell}$ below an arbitrary threshold $q_{0}=68$ are filtered out. We then concatenate all the remaining poles and residues into unique distributions used as initial parameters in the auto-differentiation approach.
We compare this combined approach to the classical SEM in Figure 8, for the same reflection coefficient as in Figures 6 and 7, i.e. the case of a slab of $\mathrm{TiO}_{2}$ In Figure 8 (a), we observe a slightly more accurate reconstruction with the combined ADC+SEM approach than with the SEM only. Let us stress that the optimization process takes, in this case, half as long with the combined approach than with just the SEM. Another significant advantage is that fewer singularities are needed with the ADC+SEM approach than with the SEM only, as shown in Figure 8 (b), where the singularities associated to both approaches are shown in a frequency window of the complex $\omega$ plane.
5 Conclusion
To conclude, we described the two approaches used in the SEMPO toolbox to retrieve complex poles, zeros and residues of arbitrary response functions. By considering the response functions of different optical systems, we showed that the improved version of the Cauchy method and the auto-differentiation based approach allow for highly accurate reconstructions of the response functions by retrieving the complex parameters of the SEM and SZF expressions.
We explained how the Accuracy-Driven Cauchy method (ADC method) improves the accuracy of the reconstructions and can be constrained to account for the stability and Hermitian-symmetry of physical systems.
We introduced an equivalent form of the SEM, namely the GDL model, which offers a different interpretation of response functions, and can still be retrieved using the auto-differentiation-based approach. This method relies on a weighted sum of loss function terms which can be used to give more importance to the average reconstruction of a target function, the maximum deviation from the target, the real part of the response function, or the imaginary part.
Although the ADC method can run much faster than the auto-differentiation approach, it is not as efficient when strong variations are observed in the response curve. As for the auto-differentiation method, it is highly dependent on the initial distributions of poles and residues which can slow down the optimization process. Consequently, we have shown how to associate the two approaches sequentially to improve the accuracy of the reconstruction, speed up the optimization process, and reduce the size of the distribution of poles.
While we have considered optical responses of photonic nanostructures to study and assess the performances of SEMPO, this method is general and can be applied to other fields of physics since it applies for any spectral response or transfer function in wave physics.
Code availability
SEMPO is available in the Zenodo repository in reference [45]